# Do We Need All the Synthetic Data? Towards Targeted Synthetic Image Augmentation via Diffusion Models

### v1

@@ Abstract
Synthetically augmenting training datasets with diffusion models has been an effective strategy for improving generalization of image classifiers. However, existing techniques struggle to ensure the diversity of generation and increase the size of the data by up to 10-30x to improve the in-distribution performance. In this work, we show that synthetically augmenting part of the data that is not learned early in training outperforms augmenting the entire dataset. By analyzing a two-layer CNN, we prove that this strategy improves generalization by promoting homogeneity in feature learning speed without amplifying noise. Our extensive experiments show that by augmenting only 30%-40% of the data, our method boosts the performance by up to 2.8% in a variety of scenarios, including training ResNet, ViT and DenseNet on CIFAR-10, CIFAR-100, and TinyImageNet, with a range of optimizers including SGD and SAM. Notably, our method applied with SGD outperforms the SOTA optimizer, SAM, on CIFAR-100 and TinyImageNet. It can also easily stack with existing weak and strong augmentation strategies to further boost the performance.
@@ 1 Introduction
Data augmentation has been essential to obtaining state-of-the-art in image classification tasks. In particular, adding synthetic images generated by diffusion models [34, 31, 35] improves the accuracy [4] and effective robustness [5] of image classification, beyond what is achieved by weak (random crop, flip, color jitter, etc) and strong data augmentation strategies (PixMix, DeepAugment, etc) [20, 21] or data augmentation using traditional generative models. [18, 7]. Existing works, however, generate synthetic images by conditioning the diffusion model on class labels [5, 4], or noisy versions of entire training data [46]. Unlike weak and strong augmentation techniques, data augmentation techniques based on diffusion models often struggle to ensure diversity and increase the size of the training data by up to 10x [4] to 30x [16] to yield satisfactory performance improvement. This raises a key question:
@@
At first, this seems implausible as adding synthetic images corresponding to only a part of the training data introduces a shift between training and test data distributions and harms the in-distribution performance. However, the recent theoretical study of Nguyen et al. [30] revealed a surprising phenomenon: when training with gradient methods, amplifying features that are not learned early in training improves the performance on the original data distribution. This is shown by comparing the training dynamics of Sharpness-Aware-Minimization (SAM) with Gradient Descent (GD). SAM is an state-of-the-art optimizer that finds flatter local minima by simultaneously minimizing the value and sharpness of the loss [15]. Empirically, [30] amplified slow-learnable features via a one-shot upsampling of examples that are not learned early in training. Theoretically, one can amplify the slow-learnable features even more to get further performance improvement. However, upsampling such examples more than once resulted in performance degradation, which contradicts the theory.
In our work, we theoretically and empirically address this contradiction. First, by analyzing a two-layer convolutional neural network (CNN), we show that SAM suppresses learning noise from the data. Our results are aligned with that of [10] identifying SAM’s ability in learning less noise as a key factor contributing to its superior generalization performance. Then, we prove that generating faithful synthetic images for the slow-learnable part of the data effectively amplifies their features without amplifying noise. In contrast, upsampling the slow-learnable part of the data amplifies their unique noise, which in turn harms the generalization performance. To generate faithful images, we use real images to guide the diffusion process. This enables synthetically augmenting the slow-learnable part of the data by up to 5x to get further performance improvement. We also prove the faster convergence of synthetic data generation over upsampling, when training with stochastic gradient methods.
@@
We conduct extensive experiments for training ResNet, ViT, and DenseNet on CIFAR10, CIFAR100 [27] and TinyImageNet [28]. We show that our synthetic data augmentation outperforms upsampling or synthetically augmenting the full dataset, and improves SGD and SAM by up to 2.8% by augmenting only 30%-40% of the data. Notably, our method applied with SGD outperforms SAM on CIFAR-100 and TinyImageNet and yields state-of-the-art performance. It can also easily stack with existing weak and strong augmentation strategies to further boost the performance.
@@ 2 Related Works
Generative Models for Augmentation. Generative models for data augmentation have become an active area of research with techniques such as GANs [45, 3, 39]. As diffusion models have emerged as a class of powerful generative models, many studies also use synthetic data from diffusion models for data augmentation. For example, Azizi et al. [4] applied diffusion models to ImageNet classification, while further studies [38, 19] explored their application in zero- or few-shot settings. Data augmentation techniques based on diffusion models often struggle to ensure diversity. Thus, recent work [23, 14, 16, 46] has focused on diversifying synthetic images to further enhance performance. Despite the promise, this line of research faces critical challenges. Most notably, these methods are often very complex, as they need to design mechanism to condition the creation of diverse prompts and optimization of conditional embeddings. This hinders their practical applications. Besides, they require to augment the dataset by up to 10x to 20x to demonstrate notable performance improvements. Generating such large volumes of data requires significant computational resources and time.
In our work, we theoretically and empirically study if this large volume of synthetic data is necessary for improving in-distribution performance. Interestingly, we show that synthetically augmenting 30%-40% of examples that are not learned early in training outperforms full data augmentation. Our method is simple and requires much lower computational resources than existing pipelines.
Sharpness-aware-minimization (SAM). SAM is an optimization technique that obtains state-of-the-art performance on a variety of tasks, by simultaneously minimizing the loss and its sharpness  [15, 44]. To do so, SAM minimizes the worst-case loss within a neighborhood around the current parameters. In doing so, it doubles the training time. SAM has also been shown to be beneficial in settings such as label noise [15, 44], out-of-distribution [37], and domain generalization [9, 40]. The superior generalization performance of SAM has been contributed to smaller Hessian spectra [15, 25, 41, 6], sparser solution [2], and benign overfitting in presence of weaker signal [10]. Most recently, SAM is shown to learn features at a more uniform speed [30]. In our work, we reveal another key property of SAM contributing to its superior performance: SAM effectively suppresses learning noise from the data. Our results suggest synthetic data generation for slower-learnable part of the data as an effective data augmentation approach for improving the generalization performance.
@@ 3 Preliminary
In this section, we introduce our theoretical framework for analyzing synthetic data augmentation with diffusion models. We base our analysis on a simplified data model that reflects the essential structure of image data, where only a small number of features are relevant to the target label. Using this model, we study the training dynamics of a two-layer convolutional neural network. We also discuss how amplifying the slow-learnable features makes training dynamics more similar to SAM.
3.1 Theoretical Setup
Data Distribution. We adopt a similar data distribution used in recent works on feature learning [1, 10, 24, 8, 26, 13, 11] to model data containing two features ${\bm{v}}_{d},{\bm{v}}_{e}$ and noise patches. Formally,
@@ 3.1 Theoretical Setup
A data point has the form $({\bm{x}},y)\sim{\mathcal{D}}(\beta_{e},\beta_{d},\alpha)\in({\mathbb{R}}^{d})^{P}\times\{\pm 1\}$, where $y\sim\text{Radamacher}(0.5)$, $0\leq\beta_{d}<\beta_{e}\in{\mathbb{R}}$, and ${\bm{x}}=({\bm{x}}^{(1)},{\bm{x}}^{(2)},\cdots,{\bm{x}}^{(P)})$ contains $P$ patches such that for some $\alpha>0$,
  • Exactly one patch is given by the fast-learnable feature $\beta_{e}\cdot y\cdot{\bm{v}}_{e}$ for some unit vector ${\bm{v}}_{e}$ with probability $\alpha$. Otherwise, the patch is given by the slow-learnable feature $\beta_{d}\cdot y\cdot{\bm{v}}_{d}$ for some unit vector ${\bm{v}}_{e}\cdot{\bm{v}}_{d}=0$.
  • The other $P-1$ patches are i.i.d. Gaussian noise $\bm{\xi}$ from ${\mathcal{N}}(0,(\sigma_{p}^{2}/d){\bm{I}}_{d})$ for some constant $\sigma_{p}$.
Here probability $\alpha$ controls the frequency of feature ${\bm{v}}_{e}$ in the data distribution. The distribution parameters $\beta_{e},\beta_{d}$ characterize the feature strength in the data. $\beta_{e}>\beta_{d}$ ensures that the fast-learnable feature is represented better in the population and thus learned faster. The faster speed of learning captures various notions of simplicity, such as simpler shape, larger magnitude, and less variation. Note that image data in practice are high-dimensional and the noises become dispersed. For simplicity, we assume $P=2$, that the noise patch is orthogonal from the two features, and that summations involving noise cross-terms $\langle\bm{\xi}_{i},\bm{\xi}_{j}\rangle$ become negligible.
@@
Empirical Risk Minimization. We consider minimizing the following empirical loss:
${\mathcal{L}}({\bm{W}})=\frac{1}{N}\sum_{i=1}^{N}\underbrace{\log(1+\exp(-y_{i}f({\bm{x}}_{i};{\bm{W}}))}_{l(y_{i}f({\bm{x}}_{i};{\bm{W}})),\text{ the logistic loss}},$ (1)
via (1) sharpness-aware minimization (SAM) [15] and (2) gradient descent (GD), whose filter-wise update rules, with some learning rate $\eta>0$, are respectively given by:
$\textbf{SAM}:{\bm{w}}_{j}^{(t+1)}$ $\!\!=\!{\bm{w}}_{j}^{(t)}\!\!-\!\eta\nabla_{{\bm{w}}_{j}^{{(t)}}}{\mathcal{L}}({\bm{W}}^{(t)}\!+\!\rho^{(t)}\nabla{\mathcal{L}}({\bm{W}}^{(t)})),\text{ where $\rho^{(t)}\!=\rho/\|\nabla{\mathcal{L}}({\bm{W}}^{(t)})\|_{F}$, $\rho>0$ },$
@@
High-level idea of SAM. By perturbing the weights with gradient ascent ($\bm{\epsilon}^{(t)}=\rho^{(t)}\nabla{\mathcal{L}}({\bm{W}}^{(t)})$), SAM looks ahead in the worst weight direction and forces the training algorithm to escape an unstable (sharp) local minimum. In practice, this has led to more generalizable solutions, while keeping track of two different gradients doubles training time and memory.
3.2 SAM Learns Features More Homogeneously
With the above setting, the alignment of ${\bm{v}}_{d},{\bm{v}}_{e}$ with weights, i.e., $\langle{\bm{w}}^{(t)},{\bm{v}}_{d}\rangle$ and $\langle{\bm{w}}^{(t)},{\bm{v}}_{d}\rangle$, indicate how much they are learned by the CNN at iteration $t$. Nguyen et al. [30] showed that SAM’s (normalized) gradient for the slow-learnable feature is larger than GD by a factor of
@@ 3.2 SAM Learns Features More Homogeneously
That is, SAM amplifies the slow-learnable feature and learns it faster than GD. In doing so, it learns features at a more homogeneous speed.
Useful (UpSample Early For Uniform Learning). Motivated by the theory, [30] proposed Useful which partitions examples to two clusters after a few training epochs and upsamples the cluster with the higher average loss once ($k=2$) and restart training on the new data distribution.
While Eq. 2 suggests that the empirical choice of $k$ should depend on the relative strength and difficulty of the features, empirically Useful with $k>2$ resulted in performance degradation, which contradicts the theory. Next, we study this discrepancy between theory and empirical results and show how synthetic data augmentation can effectively bridge the gap.
4 Learning Features More Homogeneously without Overfitting Noise
In this section, we prove that SAM suppresses learning noise from the data. Then, we show that upsampling a part of the data by a factor of $k$ encourages the model to overfit their unique noises, to an extent that scales with $k$. In contrast, synthetic data augmentation does not suffer from this issue and can effectively amplify features without overfitting noise. Finally, we verify that synthetic data augmentation has superior convergence properties over upsampling.
@@ 4.1 SAM Suppresses Learning Noise from the Data
First, we theoretically analyze how SAM suppresses learning noise in the above setting. Intuitively, as SAM pushes the learning dynamics away from sharp landscapes, it simultaneously helps the model avoid areas where certain noises concentrate. This becomes a natural defense against noise overfitting in high-curvature areas. On the other hand, gradient descent is unaware of local smoothness, so it finds solutions that may sit in a flat, noise-resilient basin.
Next, we prove that starting with the same weights ${\bm{W}}^{(t)}$, a SAM step suppresses the model’s alignment with noise directions more effectively than an equivalent gradient step, which performs the regular update. For notations, let $\bm{\Phi}$ denote the sets of noises for dataset $D$. Let ${\bm{w}}_{j,\bm{\epsilon}}$ denote the perturbed weights of filter $j$ for SAM. We then define ${\mathcal{I}}_{j,\bm{\epsilon},+}^{(t)}=\{\bm{\phi}_{i}\in\bm{\Phi}:i\in[N],\ \text{sgn}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle=\text{sgn}(y_{i})\}$ and ${\mathcal{I}}_{j,\bm{\epsilon},-}^{(t)}=\{\bm{\phi}_{i}\in\bm{\Phi}:i\in[N],\ \text{sgn}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle\neq\text{sgn}(y_{i})\}$ be the sets of noises where the sign of alignment matches or mismatches the sign of the label. We define ${\mathcal{I}}_{j,+}^{(t)}$, ${\mathcal{I}}_{j,-}^{(t)}$ accordingly for each GD weight ${\bm{w}}_{j}^{(t)}$. We measure filter-wise noise learning for a set of noises using the metric $\textbf{NoiseAlign}({\mathcal{I}},{\bm{w}}_{j}^{(t)})=\frac{1}{|{\mathcal{I}}|}\sum_{\phi\text{ in }{\mathcal{I}}}|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),\phi\rangle|$, which intuitively measures how much noise is learned by the model.
The following theorem quantifies how SAM learns noise to a smaller extent compared to GD. While we analyze the early training phase due to simplicity of analysis, at a high level, our results should still hold during the training.
@@
With sufficiently large data $N$ and model parameters $J$, and sufficiently small learning rate $\eta$ and SAM perturbation parameter $\rho$, a SAM and a GD update from the same parameters have the following property, early in training:
@@
    Inert Noises: Alignment with noises that belong to ${\mathcal{I}}_{j,-}^{(t)}$ and ${\mathcal{I}}_{j,\bm{\epsilon},-}^{(t)}$ will get closer to 00 after each update, so they will not be learned eventually by GD or SAM.
@@
    Noise Learning: The other noises will continue being learned in the sense that $|\langle{\bm{w}}_{j},{\bm{\xi}}_{i}\rangle|$ is monotonically increasing. For these noises, SAM slows down noise learning by looking ahead to noise-sensitive (sharp) directions, while GD updates “blindly”. Let $l_{i}^{(t)}=\text{sigmoid}(-y_{i}f({\bm{x}}_{i};{\bm{W}}^{(t)}))$ be the logit term. The following SAM and GD learning dynamics hold for $\bm{\xi}_{i}\in{\mathcal{I}}_{j,\bm{\epsilon},+}^{(t)}$ and $\bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{(t)}$ in terms of noise gradient:
@@
All the proof can be found in Appendix A. Our results are aligned with [10] which showed, in a different setting, that SAM’s ability to learn less noise contributes to its superior generalization.
@@
Theorem 4.1 implies that to resemble feature learning with SAM and ensure superior convergence, it is crucial to avoid magnifying noise when amplifying the slow-learnable feature. Next, we will show that synthetic data augmentation can achieve the above property, while upsampling fails to do so.
4.2 Synthetic Data Augmentation Amplifies Features But not Noise
Next, we theoretically analyze the effect of synthetically augmenting and upsampling the slow-learnable part of the data by a factor of $k$ on learning dynamics of GD. We show that while both approaches amplify the slow-learnable feature similarly, synthetic data augmentation does not magnify noise and thus has a more similar feature learning behavior to SAM. On the other hand, upsampling encourages the model to overfit the unique noises in the upsampled examples, to an extent that scales with $k$. In doing so, upsampling with $k>2$ harms the performance.

### v2

# Do We Need All the Synthetic Data? Targeted Image Augmentation via Diffusion Models

@@ Abstract
Synthetically augmenting training datasets with diffusion models has been an effective strategy for improving generalization of image classifiers. However, existing techniques struggle to ensure the diversity of generation and increase the size of the data by up to 10-30x to improve the in-distribution performance. In this work, we show that synthetically augmenting part of the data that is not learned early in training with faithful images—containing same features but different noise—outperforms augmenting the entire dataset. By analyzing a two-layer CNN, we prove that this strategy improves generalization by promoting homogeneity in feature learning speed without amplifying noise. Our extensive experiments show that by augmenting only 30%-40% of the data, our method boosts generalization by up to 2.8% in a variety of scenarios, including training ResNet, ViT, ConvNeXt, and Swin Transformer on CIFAR-10/100, and TinyImageNet, with various optimizers including SGD and SAM. Notably, our method applied with SGD outperforms the SOTA optimizer, SAM, on CIFAR-100 and TinyImageNet.
@@ 1 Introduction
Data augmentation has been essential to obtaining state-of-the-art in image classification tasks. In particular, adding synthetic images generated by diffusion models (Rombach et al., 2022; Nichol et al., 2021; Saharia et al., 2022) improves the accuracy (Azizi et al., 2023) and effective robustness (Bansal & Grover, 2023) of image classification, beyond what is achieved by weak (random crop, flip, color jitter, etc) and strong data augmentation strategies (PixMix, DeepAugment, etc) (Hendrycks et al., 2021; 2022) or data augmentation using traditional generative models (Goodfellow et al., 2020; Brock et al., 2018). Existing works, however, generate synthetic images by conditioning the diffusion model on class labels (Bansal & Grover, 2023; Azizi et al., 2023), or noisy versions of entire training data (Zhou et al., 2023). Unlike weak and strong augmentation techniques, data augmentation techniques based on diffusion models often struggle to ensure diversity and increase the size of the training data by up to 10x (Azizi et al., 2023) to 30x (Fu et al., 2024) to yield satisfactory performance improvement. This raises a key question:
@@
At first, this seems implausible as adding synthetic images corresponding to only a part of the training data introduces a shift between training and test data distributions and harms the in-distribution performance. However, recent results in the optimization literature have revealed that learning features at a more uniform speed during training improves the generalization performance (Nguyen et al., 2024). This is shown by comparing learning dynamics of Sharpness-Aware-Minimization (SAM) with Gradient Descent (GD) optimizers. SAM is a state-of-the-art optimizer that finds flatter local minima by simultaneously minimizing the value and sharpness of the loss (Foret et al., 2020). In doing so, SAM learns slow-learnable features faster than GD and achieves superior performance. This suggests that augmenting the slow-learnable part of the data to accelerate their learning can improve the generalization performance, despite slight distribution shift. Yet, how to generate such synthetic data remains an open question.
In our work, we provide a rigorous answer to the above question. First, by analyzing a two-layer convolutional neural network (CNN), we show that SAM suppresses learning noise from the data, while speeding up learning slow-learnable features. Then, we prove that generating faithful synthetic images containing slow-learnable features with different noise effectively speeds up learning such features without causing noise overfitting. To find examples with slow-learnable features, we partition the data to two parts by clustering model outputs early in training and identify the cluster with higher average loss. Then, we generate faithful images corresponding to the slow-learnable examples, by using real data to guide the diffusion process. That is, we add noise to examples that are not learned early in training and denoise them to generate faithful synthetic data (see examples in Figure 1). This enables synthetically augmenting only the slow-learnable part of the data by up to 5x to get further performance improvement. In contrast, upsampling slow-learnable examples—which appears to be a simpler and more intuitive approach—more than once could amplify the noise and significantly harm the performance, suggesting the necessity of using synthetic data. Finally, we prove the convergence properties of training on our synthetically augmented data with stochastic gradient methods.
@@
We conduct extensive experiments for training ResNet, ViT, DenseNet, and Swin Transformer on CIFAR10, CIFAR100 (Krizhevsky et al., 2009) and TinyImageNet (Le & Yang, 2015). We show that our synthetic data augmentation outperforms upsampling or synthetically augmenting the full dataset, and improves SGD and SAM by up to 2.8% by augmenting only 30%-40% of the data. Notably, our method applied with SGD outperforms SAM on CIFAR-100 and TinyImageNet and yields state-of-the-art performance. It remains effective across different diffusion models and easily stacks with existing weak and strong augmentation strategies to further boost the performance.
Figure 1: Examples of slow- and fast-learnable images and our faithful synthetic images corresponding to slow-learnable examples generated for CIFAR-10. Our synthetic data preserves features in slow-learnable images but replace noise. This amplifies slow-learnable features without magnifying noise. This is difficult to achieve with standard augmentations like random cropping or flipping, highlighting the value of generative augmentation. Additional images are given in Figure 6.
@@ 2 Related Works
Generative Models for Augmentation. There has been a recent surge of studies on synthetic data augmentation using diffusion models. For example, Azizi et al. (2023) applied diffusion models to ImageNet classification, while further studies (Trabucco et al., 2023; He et al., 2023) explored their application in zero- or few-shot settings. Despite the promise, this line of research faces critical challenges. Data augmentation techniques based on diffusion models often struggle to ensure diversity. Thus, recent works (Islam et al., 2024b; Dunlap et al., 2023; Fu et al., 2024; Zhou et al., 2023) has focused on diversifying synthetic images to further enhance performance. However, these methods are often very complex, as they need to design mechanism to condition the creation of diverse prompts and optimization of conditional embeddings. This hinders their practical applications. Besides, they require to augment the dataset by up to 10x to 30x to demonstrate notable performance improvements. Generating such large volumes of data requires significant computational resources and time.
In our work, we theoretically and empirically study if this large volume of synthetic data is necessary for improving in-distribution performance. Interestingly, we show that synthetically augmenting 30%-40% of examples that are not learned early in training outperforms full data augmentation.
Sharpness-aware-minimization (SAM). SAM is an optimization technique that obtains state-of-the-art performance on a variety of tasks, by simultaneously minimizing the loss and its sharpness  (Foret et al., 2020; Zheng et al., 2021). In doing so, it improves the generalization in expense of doubling the training time. SAM has also been shown to be beneficial in settings such as label noise (Foret et al., 2020; Zheng et al., 2021), out-of-distribution (Springer et al., 2024), and domain generalization (Cha et al., 2021; Wang et al., 2023). The superior generalization performance of SAM has been contributed to smaller Hessian spectra (Foret et al., 2020; Kaur et al., 2023; Wen et al., 2022; Bartlett et al., 2023), sparser solution (Andriushchenko & Flammarion, 2022), and benign overfitting in presence of weaker signal (Chen et al., 2022). Most recently, SAM is shown to learn features at a more uniform speed (Nguyen et al., 2024). In our work, we show that targeted synthetic data augmentation can improve generalization by making training dynamics more similar to SAM.
@@ 3 Preliminary
In this section, we introduce our theoretical framework for analyzing synthetic data augmentation with diffusion models. We also discuss SAM’s ability in learning features at a more uniform speed.
Data Distribution. We adopt a similar data distribution used in recent works on feature learning (Allen-Zhu & Li, 2020; Chen et al., 2022; Jelassi & Li, 2022; Cao et al., 2022; Kou et al., 2023; Deng et al., 2023; Chen et al., 2023) to model data containing two features ${\bm{v}}_{d},{\bm{v}}_{e}$ and noise patches.
@@
A data point has the form $({\bm{x}},y)\sim{\mathcal{D}}(\beta_{e},\beta_{d},\alpha)\in({\mathbb{R}}^{d})^{P}\times\{\pm 1\}$, where $y\sim\text{Radamacher}(0.5)$, $0\leq\beta_{d}<\beta_{e}\in{\mathbb{R}}$, and ${\bm{x}}=({\bm{x}}^{(1)},{\bm{x}}^{(2)},\cdots,{\bm{x}}^{(P)})$ contains $P$ patches.
  • Exactly one patch is given by the fast-learnable feature $\beta_{e}\cdot y\cdot{\bm{v}}_{e}$ for some unit vector ${\bm{v}}_{e}$ with probability $\alpha>0$. Otherwise, the patch is given by the slow-learnable feature $\beta_{d}\cdot y\cdot{\bm{v}}_{d}$ for some unit vector ${\bm{v}}_{e}\cdot{\bm{v}}_{d}=0$.
  • The other P1P-1 patches are i.i.d. Gaussian noise $\bm{\xi}$ from ${\mathcal{N}}(0,(\sigma_{p}^{2}/d){\bm{I}}_{d})$ for some constant $\sigma_{p}$.
Probability $\alpha$ controls the frequency of feature ${\bm{v}}_{e}$ in the data distribution. The distribution parameters $\beta_{e},\beta_{d}$ characterize the feature strength in the data. $\beta_{e}>\beta_{d}$ ensures that the fast-learnable feature is represented better in the population and thus learned faster. The faster speed of learning captures various notions of simplicity, such as simpler shape, larger magnitude, and less variation. Note that image data in practice are high-dimensional and the noises become dispersed. For simplicity, we assume $P=2$, that the noise patch is orthogonal from the two features, and that summations involving noise cross-terms $\langle\bm{\xi}_{i},\bm{\xi}_{j}\rangle$ become negligible.
@@
Empirical Risk Minimization. We consider minimizing the following empirical logistic loss:
${\mathcal{L}}({\bm{W}})=\frac{1}{N}\sum_{i=1}^{N}l(y_{i}f({\bm{x}}_{i};{\bm{W}})):=\frac{1}{N}\sum_{i=1}^{N}\log(1+\exp(-y_{i}f({\bm{x}}_{i};{\bm{W}})).$ (1)
via (1) sharpness-aware minimization (SAM) (Foret et al., 2020) and (2) gradient descent (GD), whose filter-wise update rules, with some learning rate $\eta>0$, are respectively given by:
[…]
@@
High-level idea of SAM. By perturbing the weights with gradient ascent ($\bm{\epsilon}^{(t)}=\rho^{(t)}\nabla{\mathcal{L}}({\bm{W}}^{(t)})$), SAM looks ahead in the worst weight direction and forces the training algorithm to escape an unstable (sharp) local minimum. In practice, this leads to more generalizable solutions.
SAM Learns Features More Homogeneously. With the above setting, the alignment of ${\bm{v}}_{d},{\bm{v}}_{e}$ with weights, i.e., $\langle{\bm{w}}^{(t)},{\bm{v}}_{d}\rangle$ and $\langle{\bm{w}}^{(t)},{\bm{v}}_{d}\rangle$, indicate how much they are learned by the CNN at iteration $t$. SAM’s (normalized) gradient for the slow-learnable feature is larger than GD by a factor of (Nguyen et al., 2024):
@@
That is, SAM amplifies the slow-learnable feature and learns it faster than GD. In doing so, it learns features at a more homogeneous speed. While Eq. 2 suggests that the empirical choice of $k$ should depend on the relative strength and difficulty of the features, simply upsampling examples with slow-learnable features more than once results in performance degradation, as we confirm in our experiments.
4 Learning Features Homogeneously without Overfitting Noise
In this section, we first prove that SAM suppresses learning noise from the data, while promoting homogeneous feature learning. Then, we discuss generating synthetic data to amplify features in images without magnifying their noise. This allows amplifying slow-learnable features by $k>2$ to further boost performance. Finally, we show convergence of training on our synthetically augmented data.
@@ 4.1 SAM Suppresses Learning Noise from the Data
First, we theoretically analyze how SAM suppresses learning noise in the above setting. Intuitively, as SAM pushes the learning dynamics away from sharp landscapes, it simultaneously helps the model avoid areas where certain noises concentrate. This becomes a natural defense against noise overfitting in high-curvature areas. On the other hand, gradient descent is unaware of local smoothness, so it finds solutions that may sit in a flat, noise‑resilient basin.
Fomally, we prove that starting with the same weights ${\bm{W}}^{(t)}$, a SAM step suppresses the model’s alignment with noise directions more effectively than an equivalent gradient descent step. Let $\bm{\Phi}$ denote the sets of noises for dataset $D$. Let ${\bm{w}}_{j,\bm{\epsilon}}$ denote the perturbed weights of filter $j$ for SAM. We then define ${\mathcal{I}}_{j,\bm{\epsilon},+}^{(t)}=\{\bm{\phi}_{i}\in\bm{\Phi}:i\in[N],\ \text{sgn}(\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle)=\text{sgn}(y_{i})\}$ and ${\mathcal{I}}_{j,\bm{\epsilon},-}^{(t)}=\{\bm{\phi}_{i}\in\bm{\Phi}:i\in[N],\ \text{sgn}(\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle)\neq\text{sgn}(y_{i})\}$ be the sets of noises where the sign of alignment matches or mismatches the sign of the label. We define ${\mathcal{I}}_{j,+}^{(t)}$, ${\mathcal{I}}_{j,-}^{(t)}$ accordingly for each GD weight ${\bm{w}}_{j}^{(t)}$. We measure filter-wise noise learning for a set of noises using the metric $\textbf{NoiseAlign}({\mathcal{I}},{\bm{w}}_{j}^{(t)})=\frac{1}{|{\mathcal{I}}|}\sum_{\phi\text{ in }{\mathcal{I}}}|\langle\nabla_{{\bm{w}}_{j}^{(t)}}{\mathcal{L}}({\bm{W}}^{(t)}),\phi\rangle|$, which intuitively measures how much noise is learned by the model.
The following theorem quantifies how SAM learns noise to a smaller extent compared to GD. For simplicity, we analyze the early training phase. However, our results should hold throughout the training.
@@
With controlled logit terms $l_{i}^{(t)}=\text{sigmoid}(-y_{i}f({\bm{x}}_{i};{\bm{W}}^{(t)}))$, large data size $N$, small learning rate $\eta$, and small SAM perturbation parameter $\rho$ (see Appendix A), SAM and GD updates from the same parameters have the following property, early in training:
@@
    Inert Noises: Alignment with noises that belong to ${\mathcal{I}}_{j,-}^{(t)}$ and ${\mathcal{I}}_{j,\bm{\epsilon},-}^{(t)}$ will get closer to 0 after each update, so they will not be learned eventually by GD or SAM.
@@
    Noise Learning: The other noises will continue being learned in the sense that $|\langle{\bm{w}}_{j},{\bm{\xi}}_{i}\rangle|$ is monotonically increasing. For these noises, SAM slows down noise learning by looking ahead to noise-sensitive (sharp) directions, while GD updates “blindly”. The following SAM and GD learning dynamics hold for $\bm{\xi}_{i}\in{\mathcal{I}}_{j,\bm{\epsilon},+}^{(t)}$ and $\bm{\xi}_{i}\in{\mathcal{I}}_{j,+}^{(t)}$ in terms of noise gradient:
@@
All the proof can be found in Appendix A. Our results are aligned with (Chen et al., 2023) which showed, in a different setting, that SAM can achieve benign overfitting when SGD cannot.
@@
Remark. Theorem 4.1 implies that to resemble feature learning with SAM and ensure superior convergence, it is also crucial to avoid magnifying noise when amplifying the slow-learnable feature.
4.2 Synthetic Data Augmentation To Amplify Features But not Noise
Next, we discuss finding examples containing slow-learnable features and generating synthetic data to amplify slow-learnable features without magnifying the noise in the data.
Identifying slow-learnable examples. To find slow-learnable features in the data, we find examples that are not learned robustly at the early phase of training. If an example contains at least one fast-learnable feature that is learned by the model early in training, the model relies on such features to lower the loss and potentially correctly classify the example early in training. Thus, examples that only include slow-learnable features can be identified based on loss or misclassification, or by partitioning model outputs to two clusters after a few training epochs and selecting the cluster with the higher average loss. In our experiments, we use clustering to identify examples with slow-learnable features as it yields better performance, as we confirm in our ablation studies. We note that finding examples with slow-learnable features is not the main focus of our work. Our main contribution is characterizing how to amplify slow-learnable features without amplifying noise in the data.
Amplifying slow-learnable features. Next, we show that generating synthetic data containing slow-learnable features with different noise considerably boosts the generalization performance, while upsampling slow-learnable examples amplifies the noise in the data and harms generalization.